# Climate-economy projections under shared socioeconomic pathways and net-zero scenarios

### v1

@@ 4 Conclusions
In this paper we analysed climate-economy trajectories using the cost-benefit DICE model calibrated to five baseline SSPs and six main process-based IAMs. For emission mitigation control, we considered two scenarios: achieving net-zero in 2050 and in 2100. We also considered scenarios of achieving zero industrial emission in 2050 and in 2100.
The main observation is that even if net-zero emission (or zero industrial emission) is achieved by 2050, the temperature will exceed $2^{\circ}\text{C}$ by 2100 (ranging from $2.5^{\circ}\text{C}$ to $2.7^{\circ}\text{C}$) under all five SSPs and six IAMs considered. For more lenient mitigation achieving zero emission by 2100, the temperature will be in the range between $3^{\circ}\text{C}$ and $3.7^{\circ}\text{C}$ by 2100 depending on the SSP-IAM. The social cost of carbon is raising from USD 30-50 in 2025 to USD 250-400 in 2100 depending on SSP-IAM with only very marginal impact from net-zero scenarios. This is generally consistent with other literature, see e.g. Riahi et al. (2022) and Yang et al. (2018) and provides a carbon price benchmark for policy makers. SCC does not depend significantly on the emission mitigation scenario because it measures the marginal damage to the economy from carbon emissions.
@@
Many sources of uncertainty can be considered in the DICE model, such as uncertainties in the damage function, total productivity growth, or decarbonisation function decline rate that have a significant impact on the future trajectories of climate and economy not considered in our study. These have been studied as parameter uncertainties in Nordhaus (2018). Stochastic versions of DICE where optimal policy is calculated as a decision under uncertainty solving optimal stochastic control problem are also considered in Traeger (2014); Cai and Lontzek (2019); Arandjelović et al. (2024); Shevchenko et al. (2022). Considering stochastic DICE models with exogenous functions calibrated to SSPs should also provide important benchmarks and will be considered in our future studies.

### v2

@@ 4 Conclusions
In this paper we developed climate-economy scenario trajectories using the cost-benefit DICE model calibrated to five baseline SSPs and six main process-based IAMs. For emission mitigation control, we considered two scenarios: achieving net-zero in 2050 and in 2100. We also considered scenarios of achieving zero industrial emission in 2050 and in 2100. We showed how the excess temperature-related mortality rate can be calculated for these projections, providing stress scenarios for life insurer portfolios. These scenarios of global temperature can be used for non-life insurance scenarios for events related to rising temperature. In this case, one needs to get estimates of frequency and severity of such insurable events as a function of global temperature that will directly translate to the quantifiable impact on non-life insurance products. We did not consider this and leave this to further research.
One of the main observations for the obtained climate-economy scenarios is that even if net-zero emission (or zero industrial emission) is achieved by 2050, the temperature will exceed $2^{\circ}\text{C}$ by 2100 (ranging from $2.5^{\circ}\text{C}$ to $2.7^{\circ}\text{C}$) under all five SSPs and six IAMs considered. For more lenient mitigation achieving zero emission by 2100, the temperature will be in the range between $3^{\circ}\text{C}$ and $3.7^{\circ}\text{C}$ by 2100 depending on the SSP-IAM. The social cost of carbon is raising from USD 30-50 in 2025 to USD 250-400 in 2100 depending on SSP-IAM with only very marginal impact from net-zero scenarios. This is generally consistent with other literature, see e.g. Riahi et al. (2022) and Yang et al. (2018) and provides a carbon price benchmark for policy makers. SCC does not depend significantly on the emission mitigation scenario because it measures the marginal damage to the economy from carbon emissions.
@@
Various sources of uncertainty can be considered in the DICE model, such as uncertainties in the damage function, total productivity growth, or decarbonisation function decline rate that have a significant impact on the future trajectories of climate and economy not considered in our study. These have been studied as parameter uncertainties in Nordhaus (2018). Stochastic versions of DICE where optimal policy is calculated as a decision under uncertainty solving optimal stochastic control problem are also considered in Traeger (2014); Cai and Lontzek (2019); Arandjelović et al. (2024); Shevchenko et al. (2022). Considering stochastic DICE models with exogenous functions calibrated to SSPs should also provide important benchmarks and will be considered in our future studies.
Finally, it is important to note that even though DICE model was introduced three decades ago it is regularly updated with the most recent beta release in 2023. Also, there is a recent study Folini et al. (2025) finding that the functional form of climate equations in DICE 2016 is fit for purpose, but parameters of these equations should be recalibrated. These updates should be considered in practical implementation of stress test assessment proposed in our paper.
@@ Acknowledgements
Funding Statement
This work was partially supported by JSPS KAKENHI Grant Number 21K18309.
Competing Interests
All authors declare that they have no competing interests.
Appendix A Appendix: Social cost of carbon, temperature, and concentration under various non-marker SSPs
Figure 5: Social cost of carbon ($SCC_{t}$, USD per ton of $\mathrm{CO}_{2}$), temperature ($T_{t}^{\mathrm{AT}}$), and carbon concentration ($M_{t}^{\mathrm{AT}}$) from the DICE model calibrated to baseline SSPs in AIM for scenarios of achieving net-zero emissions in 2050 (red) and 2100 (blue). Black dashed line denotes the original DICE model results.
Figure 6: Social cost of carbon ($SCC_{t}$, USD per ton of $\mathrm{CO}_{2}$), temperature ($T_{t}^{\mathrm{AT}}$), and carbon concentration ($M_{t}^{\mathrm{AT}}$) from the DICE model calibrated to baseline SSPs in GCAM for scenarios of achieving net-zero emissions in 2050 (red) and 2100 (blue). Black dashed line denotes the original DICE model results.
Figure 7: Social cost of carbon ($SCC_{t}$, USD per ton of $\mathrm{CO}_{2}$), temperature ($T_{t}^{\mathrm{AT}}$), and carbon concentration ($M_{t}^{\mathrm{AT}}$) from the DICE model calibrated to baseline SSPs in IMAGE for scenarios of achieving zero emissions in 2050 (red) and 2100 (blue). Black dashed line denotes the original DICE model results.
Figure 8: Social cost of carbon ($SCC_{t}$, USD per ton of $\mathrm{CO}_{2}$), temperature ($T_{t}^{\mathrm{AT}}$), and carbon concentration ($M_{t}^{\mathrm{AT}}$) from the DICE model calibrated to baseline SSPs in WITCH-GLOBIOM for scenarios of achieving net-zero emissions in 2050 (red) and 2100 (blue). Black dashed line denotes the original DICE model results.